Alex has 50% more money than Bob, Carl has 25% more money than Bob. All together they have a tptal of Php 3750. How much money does Alex has?
Answer by Theo(13342)

a = the amount that alex has.
b = the amount that bob has.
c = the amount that carl has.

a + b + c = 3750

a = 50% more than b, therefore a has b + .5*b = 1.5b.

c has 25% more than b, therefore c has b + .25b = 1.25b.

since a = 1.5b and c = 1.25b, then formula becomes:

1.5b + b + 1.25b = 3750

combine like terms to get 3.75b = 3750

solve for b to get b = 3750 / 3.75 = 1000.

a has 1500
b has 1000
c has 1250

total is 3750.

solution:

alex has 1500.